The area of a rectangle is five times the area of a square. If the rectangle is 50 inches long, and the width of the rectangle is the same as the length of a side of the square, find the dimensions of both the rectangle and the square.
inches (length of rectangle)
inches (width of rectangle)
inches (side length of square)

Answer by ankor@dixie-net.com(22740)

The area of a rectangle is five times the area of a square.
If the rectangle is 50 inches long, and the width of the rectangle is the same
as the length of a side of the square, find the dimensions of both the rectangle and the square.
:
Let s = side of the square, and width of the rectangle
then
s^2 = area of the square
and
50s = area of the rectangle
:
It says,"The area of a rectangle is five times the area of a square." therefore:
50s = 5s^2
divide both sides by 5s, results
10 = s
width of rect and side of square = 10 inches
:
10^2 = 100 sq/inches the area of the square
and
50*10 = 500 sq/inches the area of the rectangle (5 times the square)